# A two-circuit approach to reducing quantum resources for the quantum lattice Boltzmann method

### 20 Jan 2024 v1

@@ I Introduction
An emerging use case of quantum computers is solving computational fluid dynamics (CFD) problems [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. For example, some techniques for solving the incompressible Navier–Stokes equations rely on an efficient on a sparse linear solver. The Harrow–Hassidim–Lloyd (HHL) algorithm [2] promises an exponential speedup on such tasks, assuming the state preparation process and information readout are efficient. Lapworth [25] proposes a hybrid CFD solver based on the SIMPLE (Semi-Implicit Method for Pressure Linked Equations) algorithm and HHL. However, the number of operations HHL requires exceeds current quantum hardware capability. Instead, more proposals focus on using variational quantum algorithms (VQA) as near-term replacements for HHL to solve linear systems without any guaranteed complexity advantage. Demirdjian et al. [26] solves 1D advection-diffusion equation with Carleman linearization technique and variational quantum linear solver [27].
@@ Acknowledgements
We thank Professors J. Yepez and L. Budinski for fruitful discussions of this work. This work was supported in part by DARPA grant number HR0011-23-3-0006, the Georgia Tech Quantum Alliance, and a Georgia Tech Seed Grant.

### 11 Apr 2024 v2

@@ I Introduction
An emerging use case of quantum computers is solving computational fluid dynamics (CFD) problems [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. For example, some techniques for solving the incompressible Navier–Stokes equations rely on an efficient on a sparse linear solver. The Harrow–Hassidim–Lloyd (HHL) algorithm [2] promises an exponential speedup on such tasks, assuming the state preparation process and information readout are efficient. Lapworth [26] proposes a hybrid CFD solver based on the SIMPLE (Semi-Implicit Method for Pressure Linked Equations) algorithm and HHL. However, the number of operations HHL requires exceeds current quantum hardware capability. Instead, proposals focus on using variational quantum algorithms (VQA) as near-term replacements for HHL to solve linear systems without any guaranteed complexity advantage, which can sometimes even begin to account for nonlinaer effects [27]. For eample, Demirdjian et al. [28] solve the 1D advection-diffusion equation with a Carleman linearization technique and the variational quantum linear solver [29].
@@ Acknowledgements
We thank Professors J. Yepez and L. Budinski for fruitful discussions of this work. This work was supported in part by DARPA grant number HR0011-23-3-0006, the Georgia Tech Quantum Alliance, and a Georgia Tech Seed Grant. This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725.